OpenStudy (anonymous):

the lateral area of a cube is 576 square kilometers. what is the side length

OpenStudy (rulnick):

The lateral area is the square of the side length, times 6, since there are six faces. So we have 6x^2 = 576, where x is the side length. So x^2 = 576/6, or x^2 = 96 km^2. So take the square root to get x ...